# Convergence and error control of consistent PINNs for elliptic PDEs

### v1

@@ 10.2 Results
This solution is chosen such that it is not smooth (although its second derivatives are still continuous) to test the ability of the algorithm to handle lower regularity. The results of this experiment can be found in Table 2.
We remark that the $H^{1}$ relative errors in both experiments are reported with respect to the discrete $H^{1}(\Omega)$-norm
$\|v\|_{H^{1}(\Omega)}^{*}:=\sqrt{\left(\|v\|_{L_{2}(\Omega)}^{*}\right)^{2}+\left(\||\nabla v|\|_{L_{2}(\Omega)}^{*}\right)^{2}}$
using $500$ points in each direction.
$\widetilde{m}$ $\overline{m}$ CPINNs CPINNs PINNs PINNs
$H^{1}$ Rel error Loss ${\cal L}_{sq}^{*}$ $H^{1}$ Rel error Loss ${\cal L}_{sq}$
25 16 2.02% $2.4\cdot 10^{-5}$ 1.77% $9\cdot 10^{-6}$
100 36 0.15% $3\cdot 10^{-6}$ $0.39$% $5\cdot 10^{-6}$
225 56 0.06% $2\cdot 10^{-6}$ $0.55$% $1.9\cdot 10^{-5}$
400 76 0.07% $6\cdot 10^{-6}$ $0.58$% $2.5\cdot 10^{-5}$
625 96 0.11% $1.1\cdot 10^{-5}$ $0.48$% $2\cdot 10^{-5}$
900 116 0.07% $5\cdot 10^{-6}$ $0.34\%$ $1.3\cdot 10^{-5}$
Table 1: Results of the numerical experiments for the harmonic solution $u$. We see that the CPINNs loss function ${\cal L}_{sq}^{*}$ results in about $3-5$ times lower error compared with the original PINNs loss ${\cal L}_{sq}$.
From these results, we see that using the new loss function ${\cal L}^{*}_{sq}$ consistently gives errors that are about $3-10$ times smaller compared with the original PINNs loss. From the final loss values achieved, we conclude that this is not due to the training algorithm performing better on the new loss, but is rather due to the fact that ${\cal L}_{sq}^{*}$ gives a better control on the solution error than ${\cal L}_{sq}$. We remark that by weighting the terms in ${\cal L}_{sq}$ differently (i.e. by choosing different values of $\lambda$ in (1.7)), one may be able to improve the results for the original PINNs loss, as we have not rigorously tested this. However, our method does not require such tuning as all terms in ${\cal L}_{sq}^{*}$ should naturally carry the same weight.
We also remark that in our first experiments, the solution is very smooth. This has the consequence that the optimization error is the dominant source of error once $\widetilde{m}\geq 100$, i.e. in this regime the sampling error is already significantly smaller than the optimization error. This is why the error does not continue to decrease as the number of collocation points is increased beyond this point. By using a more effective optimization method which drives the loss ${\cal L}_{sq}^{*}$ this balance can be shifted and lower error can be obtained. On the other hand, in our second experiment, the solution $u$ is not particularly smooth. In this case, the sampling error is dominant all the way up to $\widetilde{m}=900$ collocation points and we see the error of our method decrease all the way down to this point. This is due to the fact that the optimal sampling rate decreases much more slowly because of the non-smoothness of $u$.
$\widetilde{m}$ $\overline{m}$ CPINNs CPINNs CPINNs CPINNs
$H^{1}$ Rel error Loss ${\cal L}_{sq}^{*}$ $H^{1}$ Rel error Loss ${\cal L}_{sq}$
25 16 52.4% $1.4\cdot 10^{-4}$ 46.7% $2.6\cdot 10^{-3}$
100 36 22.3% $1.4\cdot 10^{-3}$ 28.0% $6.6\cdot 10^{-3}$
225 56 9.6% $4.2\cdot 10^{-3}$ 26.7% $1.6\cdot 10^{-2}$
400 76 7.2% $5.9\cdot 10^{-3}$ 19.2% $7.7\cdot 10^{-3}$
625 96 5.5% $6.8\cdot 10^{-3}$ 25.4% $2.3\cdot 10^{-2}$
900 116 3.9% $3.5\cdot 10^{-3}$ 23.6% $2.8\cdot 10^{-2}$
Table 2: Results of the numerical experiments for the non-smooth solution $u$. We see that the CPINNs loss function ${\cal L}_{sq}^{*}$ results in about $5-10$ times lower error compared with the original PINNs loss ${\cal L}_{sq}$.
@@ 11 Concluding Remarks
We have investigated collocation methods for solving elliptic PDEs of the form (1). A collocation method seeks to generate an approximation to the true solution $u$ of (1) from point value information of the right side $f$ and the boundary value $g$. Under model class assumptions of the form $f\in{\cal F}$ and $g\in{\cal G}$ which are both unit balls of Besov spaces, we showed that there is an optimal error (called the error of optimal recovery) in recovering $u$ in the $H^{1}(\Omega)$ norm. We determined this optimal recovery error for the various model classes.
We then turned to the study of theoretical and numerical algorithms which would yield an optimal recovery. The most prominent of these methods is to minimize a certain loss function ${\cal L}$ that measures how well we have fit the data observations. We have shown that the typical loss function used in practice (based on least squares minimization) is not the correct loss and have proposed the correct loss function ${\cal L}^{*}$ defined in (7.1). We then showed that any minimizer of the correct loss ${\cal L}^{*}$ over a suitable set $\Sigma_{n}(M)$ yields a near optimal recovery of $u$.
We have then gone on to show that a modfication of PINNs which we call Consistent PINNs (CPINNs) leads to a provable near optimal recovery algorithm. The modifications needed are to use the loss ${\cal L}^{*}$ in place of the loss ${\cal L}$ usually used in PINNs. Secondly, the minimization needs to be taken over the constrained set $\Sigma_{n}(M)$ rather than all of $\Sigma_{n}$. Finally, we provide numerical experiments which demonstrate that the CPINNs loss function significantly outperforms the original PINNs loss function on elliptic PDEs.
@@
We now wish to make some further comments that one can view as serious caveats to the above theory. The first and most serious objection to the above development is that we do not provide a numerical algorithm for finding a minimizer or near minimizer of the loss ${\cal L}^{*}$ over $\Sigma_{n}(M)$. This is the same objection that can be made for almost all learning problems and certainly for PINNs. The usual numerical method used to try and solve the minimization problem in PINNs, as well as other learning problems, is to use gradient descent or some modification of gradient descent, to solve the minimization of the loss. Unfortunately, there is no proof of convergence or accuracy when applying gradient descent to non-convex losses such as ${\cal L}$ or ${\cal L}^{*}$. This has not impeded the use of gradient descent in practice with much empirical success. We have utilized such techniques as well in our numerical examples.
@@
We do not want to enter into a discussion of the art of using gradient descent in optimization. However, we want to address two issues that arise relevant to this paper. The first of these is the fact that we must optimize the loss over $\Sigma_{n}(M)$ rather than $\Sigma_{n}$ itself to gain the proof of optimality. This is a serious numerical burden. We have ignored this issue in our numerical experiments much like one ignores the lack of a provable gradient descent algorithm. One could implement the restricted approximation term by adding a penalty to the loss ${\cal L}^{*}$ to guarantee that at each step of the optimization one remains in the constrained set $\Sigma_{n}(M)$ (see [4]) but this does not ease the computational burden and one is still faced with a lack of convergence of the numerical optimization.
@@
Finally, let us point out another issue which is understood by practitioners but not usually pointed out in the theory. Neural network spaces are a very unstable manifold. By this we mean that changing parameters very slightly can have a huge effect on the output of the neural network. That is, the mapping from parameters to output in neural networks is very unstable. This instability appears more in deep networks than in shallow networks. Avoiding this lack of stability is an art in practice that manifests itself in judiciously choosing the starting parameters and the learning rates (step size in gradient descent). The theoretical aspects of this lack of stability, and how to avoid the instability theoretically, is addressed in a series of recent papers [8, 29, 28, 30]. There remains a serious gap between theoretical algorithms for learning and their numerical implementation.
@@ 12.1 Local approximation by polynomials
It is useful to use the following modified form of Whitney’s theorem. For any cube $I\subset\mathbb{R}^{d}$, any $0<p<\infty$, and $f\in L_{p}(I)$, we define
@@
where $I_{r{\rm\bf h}}:=\{{\rm\bf x}:[{\rm\bf x},{\rm\bf x}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}r}{\rm\bf h}]\subset I\}$. This is called the averaged modulus of smoothness of $f$. It is known that $w_{r}$ is equivalent to $\widetilde{w}_{r}$ (see §2 of [18]),
@@ 12.3 Besov spaces and piecewise polynomial approximation
We now recall how membership in $B_{p}^{s}(\Omega)$ guarantees a rate of approximation by piecewise polynomials. In actuality the membership of a function in a Besov space can be characterized by its rate of approximation by piecewise polynomials. However, we only need the results that prove that a function in a Besov space can be approximated by piecewise polynomials with a certain accuracy and therefore only concentrate on proving results of this type.
For each $k\geq 0$, we define ${\cal D}_{k}$ to be the partition of $\Omega$ into dyadic cubes $I$ of side length $2^{-k}$. Here we take a dyadic cube $I\in{\cal D}_{k}$ to be the tensor product of the dyadic intervals $[(j-1)2^{-k},j2^{-k})$, $1\leq j\leq 2^{k}$. We then define
@@
to be the space of all piecewise polynomials of order $r$ that are subordinate to the partition ${\cal D}_{k}$. In other words, a function $S\in{\cal S}_{k}$ if an only the restriction of $S$ on each $I\in{\cal D}_{k}$ belongs to ${\cal P}_{r}$.
@@
where $\chi_{I}$ is the characteristic function of $I$. Notice that $S_{k}(x)$ is defined pointwise for each $x\in\Omega$, We extend $S_{k}$ to all of $\overline{\Omega}$ by continuity.
@@
$\|f-S_{k}\|^{p}_{L_{p}(\Omega)}\leq C^{p}\sum_{I\in{\cal D}_{k}}\widetilde{w}_{r}(f,2^{-k})^{p}_{L_{p}(I)}\leq C^{p}\widetilde{w}_{r}(f,2^{-k})^{p}_{L_{p}(\Omega)}\leq C^{p}|f|^{p}_{B_{p}^{s}(\Omega)}2^{-ksp},$ (12.16)
where we used the subadditivity (12.6) and the fact that ${\cal D}_{k}$ is a partition of $\Omega$. This proves (12.14) in the case $p<\infty$. When $p=\infty$, these inequalities follow directly from (12.15) and the fact that $\|\cdot\|_{L_{\infty}(I)}\leq\|\cdot\|_{L_{\infty}(\Omega)}$ for each $I\subset\Omega$. ∎
@@ 12.4 Piecewise polynomial approximation in $L_{\tau}(\Omega)$
$\|R_{k}\|_{L_{p}(\Omega)}\leq C[\|f-S_{k}\|_{L_{p}(\Omega)}+\|f-S_{k-1}\|_{L_{p}(\Omega)}]\leq C|f|_{B_{p}^{s}(\Omega)}2^{-ks},\quad k\geq 1,$ (12.19)
@@
Case 2: $\tau<\infty$. Similarly to (12.21) and using the comparison of polynomial norms of (12.11), we find that
@@
$\leq$ $C^{\tau}2^{kd(\tau/p-1)}\left[\sum_{I\in{\cal D}_{k}}\widetilde{w}_{r}(f,2^{-k+1})_{L_{p}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\overline{I}})}^{p}\right]^{\tau/p}\leq 2^{d\tau/p}C^{\tau}2^{kd(\tau/p-1)}\left[\sum_{\overline{I}\in{\cal D}_{k-1}}\widetilde{w}_{r}(f,2^{-k+1})_{L_{p}(\overline{I})}^{p}\right]^{\tau/p}$
@@ 12.6 The proofs of the Theorems on optimal recovery
12.6.3 Proof of Theorem 3.3
We first prove the right inequality of (3.21). Given $m\geq 1$, let ${\cal X}$ be a collection of $\widetilde{m}:=\lfloor{\frac{m}{2}\rfloor}$ data sites in $\overline{\Omega}$ for optimal sampling of the functions in ${\cal F}$. Similarly let ${\cal Z}$ be a collection of $\overline{m}:=\lfloor{\frac{m}{2}\rfloor}$ data sites in $\partial\Omega$ for optimal sampling of the functions in ${\cal G}$. Given $u\in{\cal U}$ let $f=-\Delta u$ and $g$ be the trace of $u$ on $\partial\Omega$. We know that there is an $\widehat{f}\in H^{-1}(\Omega)$ which is an optimal $H^{-1}(\Omega)$ recovery of $f$ using the point samples at ${\cal X}$. That is, we have $\|\widetilde{f}-\widehat{f}\|_{H^{-1}(\Omega)}\leq Cm^{-\alpha}$ whenever $\widetilde{f}\in{\cal F}$ shares the same data as $f$. Similarly, let $\widehat{g}$ be an optimal $H^{1/2}(\partial\Omega)$ recovery of $g$ using the data at ${\cal Z}$. Then, we have $\|\widetilde{g}-\widehat{g}\|_{H^{1/2}(\partial\Omega)}\leq Cm^{-\beta}$ whenever $\widetilde{g}$ is a function in ${\cal G}$ which shares the same data as $g$. Now, define $\widehat{u}$ as the solution to (1) with right side $\widehat{f}$ and boundary value $\widehat{g}$. Let $\widetilde{u}$ be any function from ${\cal U}$ which shares the same data as $u$. From (1.3), we have
$\|\widetilde{u}-\widehat{u}\|_{H^{1}(\Omega)}\leq Cm^{-\min(\alpha,\beta)},$
where $C$ is independent of $m$ and the data. Since this holds for all $u\in{\cal U}$, this proves the upper inequality in (3.21)
To prove the lower inequality we consider two cases. First suppose that $\alpha\leq\beta$. Given any set ${\cal X}$ of $m$ data sites in $\overline{\Omega}$ at which we sample the $f\in{\cal F}$, we can find two functions $f_{1},f_{2}\in{\cal F}$ that have the same values on ${\cal X}$ and yet $\|f_{1}-f_{2}\|_{H^{-1}(\Omega)}\geq cm^{-\alpha}$ where $c$ does not depend on $m$. For $i=1,2$, we let $u_{i}$ satisfy (1) with right side $f_{i}$ and with $g_{i}\equiv 0$. Then, (1.3) says that
$\|u_{1}-u_{2}\|_{H^{1}(\Omega)}\geq cm^{-\alpha}=cm^{-\min(\alpha,\beta)}.$
This proves the lower inequality in (3.21) in this case. A similar argument applies when $\beta<\alpha$. $\Box$

### v2

@@ 10.2 Results
The solution $u$ vanishes on the boundary (i.e. $g=0$) and is much less smooth than the preceding two experiments. For this reason, we use a significantly larger network ($L=3$, $W=15$), more collocation points ($10$, $20$, $30$ and $40$ in each direction), and more training iterations ($1000$ steps). The results are presented in Table 3.
$\widetilde{m}$ $\overline{m}$ ${\cal L}_{sq,1}$ ${\cal L}_{sq,1}$ ${\cal L}_{sq,\lambda(\overline{m})}$ ${\cal L}_{sq,\lambda(\overline{m})}$ ${\cal L}^{*}_{sq,\gamma}$ ${\cal L}^{*}_{sq,\gamma}$ ${\cal L}^{*}_{sq,2}$ ${\cal L}^{*}_{sq,2}$
Rel Error Loss Rel error Loss Rel error Loss Rel error Loss
100 36 $8.4\%$ $9.0\cdot 10^{-4}$ $3.8\%$ $3.3\cdot 10^{-3}$ $10.4\%$ $1.0\cdot 10^{-2}$ $18.9\%$ $6.1\cdot 10^{-3}$
400 76 $18.8\%$ $5.2\cdot 10^{-2}$ $1.8\%$ $4.4\cdot 10^{-2}$ $5.1\%$ $2.2\cdot 10^{-2}$ $5.3\%$ $2.9\cdot 10^{-2}$
900 116 $10.9\%$ $1.1\cdot 10^{-2}$ $2.4\%$ $3.9\cdot 10^{-2}$ $3.1\%$ $1.9\cdot 10^{-2}$ $2.5\%$ $2.1\cdot 10^{-2}$
1600 156 $6.1\%$ $6.3\cdot 10^{-3}$ $1.2\%$ $7.5\cdot 10^{-3}$ $11.1\%$ $6.4\cdot 10^{-2}$ $1.4\%$ $1.1\cdot 10^{-2}$
Table 3: Experimental results on the non-smooth problem. All relative errors are in the $H^{1}(\Omega)$-norm, computed using a grid with $500$ points in each direction. We see that the properly weighted PINNs loss, ${\cal L}_{sq,\lambda(\overline{m})}$, and the consistent PINNs with $L_{2}$-penalty on the domain, ${\cal L}^{*}_{sq,2}$, perform significantly better than both other loss functions. Interestingly, the performance of the consistent loss ${\cal L}^{*}_{sq,\gamma}$ deteriorates significantly with the largest number of collocation points.
Our numerical experiments seem to imply that the optimization method plays a critical role in the final accuracy achieved when using any of these loss functions. Indeed, there is a fairly wide variation in the computed final loss values. Accounting for this, in the first two tests, it seems that the correctly weighted PINNs loss ${\cal L}_{sq,\lambda(\overline{m})}$ performs the best. The loss (7.2) with $L_{2}$-penalty on the residual, ${\cal L}^{*}_{sq,2}$, is performing similarly well. However, more experiments are needed to confirm our observations.
The results of Experiment 3 are more conclusive. The solution in this case is much less smooth and more difficult to fit using neural networks than in the first two experiments. In this experiment, we observe that optimizing the correctly weighted PINNs loss ${\cal L}_{sq,\lambda(\overline{m})}$ and the loss (7.2) with $L_{2}$-penalty on the domain, ${\cal L}^{*}_{sq,2}$, result in significantly lower solution error than the other two loss functions. Interestingly, the loss ${\cal L}^{*}_{sq,\gamma}$ performs nearly as well up until the final experiment where the collocation grid has $40$ points in each direction.
Our interpretation of these results is the following. The (unweighted) original PINNs loss does not adequately fit the boundary values, while the loss (7.2) results in a less smooth numerical solution due to the optimization of the non-smooth $L_{\gamma}$-norm. Both the losses ${\cal L}_{sq,\lambda(\overline{m})}$ and ${\cal L}^{*}_{sq,2}$ overcome these issues by correctly weighting the boundary error and also resulting in neural network solutions which are smooth. This is corroborated by the solution plots in each case, which are shown in Figures 10.2 and 10.3. Observe that the solutions minimizing both of the losses ${\cal L}_{sq,1}$ and ${\cal L}^{*}_{sq,\gamma}$ fail to accurately fit the boundary values. However, the solution for ${\cal L}_{sq,1}$ is relatively smooth, while the solution for ${\cal L}^{*}_{sq,\gamma}$ is significantly less smooth (especially near the boundary). The solutions for both ${\cal L}_{sq,\lambda(\overline{m})}$ and ${\cal L}^{*}_{sq,2}$ are fairly smooth and fit the boundary values well.
Figure 10.2: Experiment 3: Solution plots for the (unweighted) original PINNs loss ${\cal L}_{sq,1}$ (left) and the consistent PINNs loss ${\cal L}^{*}_{sq,\gamma}$ (right). We see that both fail to accurately fit the boundary values. Note that the consistent PINNs solution is much less smooth near the boundary.
Figure 10.3: Experiment 3: Solution plots for the correctly weighted PINNs loss ${\cal L}_{sq,\lambda(\overline{m})}$ (left) and the consistent PINNs loss with $L_{2}$-penalty on the domain ${\cal L}^{*}_{sq,2}$ (right). We see that both accurately fit the boundary values and remain fairly smooth.
Based upon our numerical tests, we recommend using either the correctly weighted PINNs loss ${\cal L}_{sq,\lambda(\overline{m})}$, see (8.5), or the loss (7.2) with $L_{2}$-penalty on the domain, ${\cal L}^{*}_{sq,2}$, especially for difficult problems with limited regularity. However, our observations are naturally quite limited and significant additional experiments are needed to draw stronger conclusions.
@@ 11 Concluding Remarks
We have investigated collocation methods for solving elliptic PDEs of the form (1). Under model class assumptions of the form $f\in{\cal F}$ and $g\in{\cal G}$, which are both unit balls of Besov spaces, we showed that there is an optimal error (called the error of optimal recovery) in recovering $u$ in the $H^{1}(\Omega)$ norm. We determined this optimal recovery error for the various model classes. Our theoretical analysis was performed for $\Omega=(0,1)^{d}$, but it is well-known that the results carry over to any Lipschitz domain (see [58]).
We then turned to the study of theoretical algorithms which would yield an optimal recovery. The most prominent of these methods is to minimize a certain loss function ${\cal L}$ that measures how well we have fit the data observations. While the typical loss function used in PINNs does not provide an a posteriori bound on the error, we have introduced a variety of modified loss functions which do provide such a bound. We call the use of any one of these theoretically justified loss functions consistent PINNs. Finally, we perform several numerical experiments which compare these different loss functions and show their favorable performance in comparison to the typical loss used in PINNs.
@@
We now wish to make some further comments that one can view as serious caveats to the above theory. The first and most serious objection to the above development is that we do not provide a numerical algorithm for finding a minimizer or near minimizer of the suggested losses over $\Sigma_{n}(M)$. This is the same objection that can be made for almost all learning problems and certainly for PINNs. The usual approach in PINNs is to use gradient descent or some modification of gradient descent to solve the minimization of the loss. Unfortunately, there is no proof of convergence or accuracy when applying gradient descent to non-convex losses such as ${\cal L}$ or ${\cal L}^{*}$. This has not impeded the use of gradient descent in practice with much empirical success. We have utilized such techniques as well in our numerical examples.
@@
We do not enter into a discussion of the art of using gradient descent in optimization. However, we want to address two issues that are relevant to this paper. The first is the fact that we must optimize the loss over $\Sigma_{n}(M)$ rather than $\Sigma_{n}$ itself to gain the proof of optimality. This is a serious numerical burden. We have ignored this issue in our numerical experiments much like one ignores the lack of a provable gradient descent algorithm. One could implement the restricted approximation by adding a penalty to the consistent losses to guarantee that at each step of the optimization one remains in the constrained set $\Sigma_{n}(M)$ (see [5]) but this does not ease the computational burden and one is still faced with a lack of convergence of the numerical optimization.
@@
The second issue is understood by practitioners but not usually pointed out in the theory. Neural network spaces are a very unstable manifold in the sense that changing parameters very slightly can have a huge effect on the output of the neural network. That is, the mapping from parameters to output in neural networks is very unstable. This instability appears more in deep networks than in shallow networks. Avoiding this lack of stability is an art in practice that manifests itself in judiciously choosing the starting parameters and the learning rates (step size in gradient descent). The theoretical aspects of this lack of stability, and how to avoid the instability theoretically, is addressed in a series of recent papers [9, 36, 35, 37]. However, there remains a serious gap between theoretical algorithms for learning and their numerical implementation.
@@ 12.1 Local approximation by polynomials
We also will use the following modified form of Whitney’s theorem. For any cube $I\subset\mathbb{R}^{d}$, any $0<p<\infty$, and $f\in L_{p}(I)$, we define
@@
where $I_{r{\rm\bf h}}:=\{{\rm\bf x}:[{\rm\bf x},{\rm\bf x}+{r}{\rm\bf h}]\subset I\}$. This is called the averaged modulus of smoothness of $f$. It is known that $w_{r}$ is equivalent to $\widetilde{w}_{r}$ (see §2 of [23]),
@@ 12.3 Besov spaces and piecewise polynomial approximation
We now recall how membership in $B_{p}^{s}(\Omega)$ guarantees a rate of approximation by piecewise polynomials. In actuality, the membership of a function in a Besov space can be characterized by its rate of approximation by piecewise polynomials. However, we only need the results that prove that a function in a Besov space can be approximated by piecewise polynomials with a certain accuracy and therefore only concentrate on proving results of this type.
Recall that for $k\geq 0$, we denoted by ${\cal D}_{k}$ the partition of $\Omega$ into dyadic cubes $I$ of side length $2^{-k}$, where the dyadic cube $I\in{\cal D}_{k}$ is the tensor product of the dyadic intervals $[(j-1)2^{-k},j2^{-k})$, $1\leq j\leq 2^{k}$. We then define
@@
to be the space of all piecewise polynomials of order $r$ that are subordinate to the partition ${\cal D}_{k}$. In other words, a function $S\in{\cal S}_{k}$ if and only if the restriction of $S$ on each $I\in{\cal D}_{k}$ belongs to ${\cal P}_{r}$.
@@
where $\chi_{I}$ is the characteristic function of $I$. Notice that $S_{k}(x)$ is defined pointwise for each $x\in\Omega$. We extend $S_{k}$ to all of $\overline{\Omega}$ by continuity.
@@
$\|f-S_{k}\|^{p}_{L_{p}(\Omega)}\lesssim\sum_{I\in{\cal D}_{k}}\widetilde{w}_{r}(f,2^{-k})^{p}_{L_{p}(I)}\lesssim\widetilde{w}_{r}(f,2^{-k})^{p}_{L_{p}(\Omega)}\lesssim w_{r}(f,2^{-k})^{p}_{L_{p}(\Omega)}\lesssim|f|^{p}_{B_{p}^{s}(\Omega)}2^{-ksp},$ (12.16)
where we used the subadditivity (12.6) and the fact that ${\cal D}_{k}$ is a partition of $\Omega$. Here we have also used the equivalence (12.4) of $\widetilde{w}_{r}$ and $w_{r}$ and (2.8). This proves (12.14) in the case $p<\infty$. When $p=\infty$, these inequalities follow directly from (12.15) and the fact that $\|\cdot\|_{L_{\infty}(I)}\leq\|\cdot\|_{L_{\infty}(\Omega)}$ for each $I\subset\Omega$. ∎
@@ 12.4 Piecewise polynomial approximation in $L_{\tau}(\Omega)$
$\|R_{k}\|_{L_{p}(\Omega)}\leq\|f-S_{k}\|_{L_{p}(\Omega)}+\|f-S_{k-1}\|_{L_{p}(\Omega)}\leq C|f|_{B_{p}^{s}(\Omega)}2^{-ks},\quad k\geq 1,$ (12.19)
@@
Case 2: $p\leq\tau<\infty$. Similarly to (12.21) and using the comparison of polynomial norms of (12.11), we find that
@@
$\leq$ $C^{\tau}2^{kd(\tau/p-1)}\left[\sum_{I\in{\cal D}_{k}}\widetilde{w}_{r}(f,2^{-k+1})_{L_{p}({\overline{I}})}^{p}\right]^{\tau/p}\leq 2^{d\tau/p}C^{\tau}2^{kd(\tau/p-1)}\left[\sum_{\overline{I}\in{\cal D}_{k-1}}\widetilde{w}_{r}(f,2^{-k+1})_{L_{p}(\overline{I})}^{p}\right]^{\tau/p}$